What is an equation of the line that passes through the point (−2,−4) and is perpendicular to the line x+5y=15. Express your answer in slope-intercept form.
y=mx+b
x+15=5y
15x/5
y=3

Answer by mananth(16946)

1 x + 5 y = 15
Find the slope of this line

5 y = -1 x + 15
Divide by 5
y = -1/5 x + 3
Compare this equation with y=mx+b
slope m = - 1/5

The slope of a line perpendicular to the above line will be the negative reciprocal =5
m1*m2=-1


m= 5 ,point ( -2 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = -10.00 + b
b= 6
m= 5
The required equation is y = 5 x + 6
m.ananth@hotmail.ca